How much coffee costing 85 cents a pound must be added to 150 pounds costing 90 cents a pound to make a mixture costing 87 cents a pound?
Answer by stanbon(75887)

How much coffee costing 85 cents a pound must be added to 150 pounds costing 90 cents a pound to make a mixture costing 87 cents a pound?
-------------------------
value + value = value
85x + 90*150 = 87(x+150)
----
85x + 90*150 = 87x + 87*150
2x = 3*150
x = 3*75
x = 235 lbs. (amt of 85 cent coffee needed in the mixture)
=============================================================
Cheers,
Stan H.